Express the following surfaces in spherical coordinates :


yz=2


1
Expert's answer
2022-01-22T15:23:22-0500

Solution:

"yz=2\ ...(i)"

We know that

"y=r\sinθ\sinϕ \\
z=r\cos{\theta}"

Putting these values in (i),

"yz= r\sinθ\sinϕ\ r\cos{\theta}
\\=r^2\sin{\theta}\cos{\theta}\cos{\phi}
\\=\frac{1}{2}r^2\sin{2\theta}\cos{\phi}"

Now,

"yz=2
\\ \Rightarrow \frac{1}{2}r^2\sin{2\theta}\cos{\phi}=2
\\ \Rightarrow r^2\sin{2\theta}\cos{\phi}=4, \ r>0,\ \theta\in[0,\pi], \ \phi\in[0,2\pi]"